Thank you for getting back to me so soon.
I have two more that I would like to check my answers.
81.72 + 36.067 + -48.93 = my answer is 68.857 is this correct?
-3.69 - (-1.527) = -2.163,is this correct as well?

Answer by jim_thompson5910(35256)

81.72 + 36.067 + -48.93 = 117.787 + -48.93 = 68.857


So you are correct.

--------------------------


-3.69 - (-1.527) = -3.69 + 1.527 = -2.163


Once again, you are correct.